# MVG-CRPS: A Robust Loss Function for Multivariate Probabilistic Forecasting

### v1

@@ 3.2 MVG-CRPS as Loss Function for Multivariate Forecasting
In multivariate probabilistic forecasting, proper scoring rules such as the log-score (Eq. (4)) and the energy score ($\operatorname{ES}$) are used to evaluate predictive performance. The energy score generalizes CRPS to assess probabilistic forecasts of vector-valued random variables (Gneiting and Raftery,, 2007):
@@
$\operatorname{ES}(F,\mathbf{z})=\frac{1}{n}\sum_{i=1}^{n}\lVert\boldsymbol{Z}_{i}-\mathbf{z}\rVert^{\beta}-\frac{1}{2n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\lVert\boldsymbol{Z}_{i}-\boldsymbol{Z}^{\prime}_{j}\rVert^{\beta},$ (11)
@@
where $\boldsymbol{S}_{t}=\operatorname{diag}(\{\lambda_{t}^{1},\ldots,\lambda_{t}^{N}\})$ is a diagonal matrix containing the eigenvalues $\lambda_{t}^{i}$ of $\boldsymbol{\Sigma}_{t}$, and $\boldsymbol{U}_{t}$ is the orthonormal matrix of corresponding eigenvectors. We then define
@@
where $\mathbf{w}_{t}$ is a random vector with each element following a standard Gaussian distribution. We can then apply Eq. (8) individually to each element and obtain a new loss function mimicking Eq. (9) for multivariate problem:
$\mathcal{L}=\sum_{t=1}^{T}\sum_{i=1}^{N}\sqrt{\lambda_{t}^{i}}\operatorname{CRPS}\left(F,{w}_{i,t}\right).$ (15)
The primary advantage of this loss function lies in its ability to leverage the closed-form expression of $\operatorname{CRPS}$ by decorrelating the time series variables through whitening. While a potential limitation is that $\operatorname{CRPS}$, when applied to each transformed (whitened) variable, primarily evaluates the marginal distributions in the transformed space, the whitening process itself is derived from the original covariance matrix. As a result, the optimization process can still capture the covariance structure of the original distribution. Consequently, the loss function implicitly addresses certain aspects of the covariance, allowing $\operatorname{CRPS}$ to retain indirect sensitivity to the joint behavior of the original variables.
@@ 4.1 Datasets and Models
We apply the proposed MVG-CRPS to two multivariate forecasting tasks: multivariate autoregressive forecasting and univariate Seq2Seq forecasting. For the first task, we employ two benchmark models: the RNN-based GPVar (Salinas et al.,, 2019) and a decoder-only Transformer (Radford et al.,, 2018). For the second task, we use the N-HiTS model, which is based on multi-layer perceptrons (MLPs) (Challu et al.,, 2023).
@@
To generate the distribution parameters for probabilistic forecasting, we employ a Gaussian distribution head based on the hidden state $\mathbf{h}_{i,t}$ produced by the model. Specifically, for the multivariate autoregressive forecasting, following Salinas et al., (2019), we parameterize the mean vector as $\boldsymbol{\mu}\left(\mathbf{h}_{t}\right)=\left[\mu_{1}\left(\mathbf{h}_{1,t}\right),\dots,\mu_{N}\left(\mathbf{h}_{N,t}\right)\right]^{\top}\in\mathbb{R}^{N}$ and adopt a low-rank-plus-diagonal parameterization of the covariance matrix $\boldsymbol{\Sigma}\left(\mathbf{h}_{t}\right)=\operatorname{diag}{(\mathbf{d}_{t})}+\boldsymbol{L}_{t}\boldsymbol{L}_{t}^{\top}$, where $\mathbf{d}_{t}=[d_{1}\left(\mathbf{h}_{1,t}\right),\ldots,d_{N}\left(\mathbf{h}_{N,t}\right)]^{\top}\in\mathbb{R}_{+}^{N}$ and $\boldsymbol{L}_{t}=\left[\mathbf{l}_{1}\left(\mathbf{h}_{1,t}\right),\dots,\mathbf{l}_{N}\left(\mathbf{h}_{N,t}\right)\right]^{\top}\in\mathbb{R}^{N\times R}$, $R\ll N$ is the rank parameter. Here, $\mu_{i}(\cdot)$, $d_{i}(\cdot)$, and $\mathbf{l}_{i}(\cdot)$ are the mapping functions that generate the mean and covariance parameters for each time series $i$ based on the hidden state $\mathbf{h}_{i={1:N},t}$. In practice, we use shared mapping functions across all time series, denoted as $\mu_{i}=\tilde{\mu}$, $d_{i}=\tilde{d}$, and $\mathbf{l}_{i}=\tilde{\mathbf{l}}$. This approach enables the use of random subsets of time series (i.e., batch size $B\leq N$) for model optimization in each iteration, making it feasible to apply our method to high-dimensional time series datasets. Similarly, in the univariate Seq2Seq forecasting task, the mean $\boldsymbol{\mu}\left(\mathbf{h}_{i}\right)$ and covariance $\boldsymbol{\Sigma}(\mathbf{h}_{i})$ are defined over the forecast horizon for each specific time series, based on the hidden states $\mathbf{h}_{i,t=T+1:T+Q}$. As a result, we can model the joint distribution $p\left(\mathbf{z}_{i,{T+1}:{T+Q}}\right)$ over the forecasted values.
@@
We implemented our models using PyTorch Forecasting (Beitner,, 2020), with input data consisting of lagged time series values and covariates. We conducted extensive experiments on diverse real-world time series datasets obtained from GluonTS (Alexandrov et al.,, 2020). The prediction horizon ($Q$) and the number of rolling evaluations were taken from the configuration within GluonTS, where we follow the default setting by equating the context range to the prediction range, i.e., $P=Q$. Each dataset was sequentially split into training, validation, and testing sets. Each time series was individually normalized using a scaler fitted to its own training data. Predictions were then rescaled to their original values for computing evaluation metrics. For comprehensive details on the experimental setup, we refer readers to the Supplementary Materials (SM).
@@ 4.2 Quantitative Evaluation
We evaluate the proposed MVG-CRPS loss by comparing it with models trained using the log-score and energy score. The evaluation metrics are $\operatorname{CRPS}_{\text{sum}}$ and energy score (Eq. (11)). $\operatorname{CRPS}_{\text{sum}}$ is computed by summing both the forecast and ground-truth values across all time series and then calculating the $\operatorname{CRPS}$ (Eq. (7)) over the resulting sums (Salinas et al.,, 2019; Drouin et al.,, 2022; Ashok et al.,, 2024):
$\operatorname{\mathbb{E}}_{t}\left[\operatorname{CRPS}\left(F_{t},\sum_{i}z_{i,t}\right)\right],$ (16)
where the empirical $F_{t}$ is calculated by aggregating samples across time series, with 100 samples drawn to compute $\operatorname{CRPS}_{\text{sum}}$. The evaluation using the energy score is provided in the SM.
Table 1 presents a comparison of $\operatorname{CRPS}_{\text{sum}}$ across different scoring rules in the multivariate autoregressive forecasting task, underscoring the effectiveness of the proposed MVG-CRPS loss. MVG-CRPS consistently outperforms the log-score across most datasets. For instance, on the $\mathtt{electricity}$ and $\mathtt{tourism}$ datasets, MVG-CRPS achieves scores of 0.0249 and 0.2004, respectively, outperforming the scores of 0.0419 and 0.2217 achieved by the log-score. This indicates that MVG-CRPS leads to models that produce higher-quality forecasts. As discussed in later sections, this improvement is attributed to MVG-CRPS being less sensitive to outliers.
Table 1: Comparison of $\operatorname{CRPS}_{\text{sum}}$ across different scoring rules in the multivariate autoregressive forecasting task. The best scores are in boldface. Scores for MVG-CRPS are underlined when they are not the best overall but are better than the log-score. Mean and standard deviation are obtained from 10 runs of each model.
@@
Table 2: Comparison of training cost (in minutes) on GPVar across different scoring rules in the multivariate autoregressive forecasting task.
@@ 4.3.1 Multivariate Autoregressive Forecasting
Figure 2: Comparison of the output covariance matrices $\boldsymbol{\Sigma}_{t}$ produced by GPVar on the $\mathtt{elec\_weekly}$ dataset. Top row and bottom row show covariance matrices from the model trained using the log-score and MVG-CRPS, respectively. For visual clarity, covariance values are clipped between 0 and 0.6.
Figure 3: Comparison of probabilistic forecasts produced by GPVar on the $\mathtt{electricity}$ dataset. The top row shows the forecasts from the model trained using the log-score, while the bottom row from the model trained using MVG-CRPS.
@@ 4.3.2 Univariate Seq2Seq Forecasting
Another example is illustrated in Fig. 5, which showcases the probabilistic forecasts generated by the N-HiTS model on the $\mathtt{m4\_hourly}$ dataset. We observe notable differences in both the sharpness and calibration of the forecasts. Specifically, the model trained with MVG-CRPS produces slightly narrower prediction intervals, indicating greater confidence in its predictions compared to the log-score-trained model. In contrast, the log-score-trained model produces unnecessarily wide prediction intervals for simple time series with strong cyclical patterns. Additionally, the MVG-CRPS-trained model demonstrates higher accuracy in longer-term forecasts.
These qualitative evaluations underscore the advantages of using MVG-CRPS as a loss function for training probabilistic forecasting models. By reducing sensitivity to outliers and providing more stable covariance estimates, MVG-CRPS enhances the model to generate accurate and reliable probabilistic forecasts.
Figure 4: Comparison of the output covariance matrices $\boldsymbol{\Sigma}_{i}$ produced by N-HiTS on the $\mathtt{traffic}$ dataset. For visual clarity, covariance values are clipped between 0 and 1.0.
Figure 5: Comparison of probabilistic forecasts produced by N-HiTS on the $\mathtt{m4\_hourly}$ dataset.

### v2

@@ 3.2 MVG-CRPS as Loss Function for Multivariate Forecasting
In multivariate probabilistic forecasting, proper scoring rules such as the log-score (Eq. (4)) and the energy score ($\operatorname{ES}$) are used to evaluate predictive performance. The energy score generalizes CRPS to assess probabilistic forecasts of vector-valued random variables (Gneiting & Raftery, 2007):
@@
$\operatorname{ES}(F,\mathbf{z})=\frac{1}{n}\sum_{i=1}^{n}\lVert\boldsymbol{Z}_{i}-\mathbf{z}\rVert^{\beta}-\frac{1}{2n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\lVert\boldsymbol{Z}_{i}-\boldsymbol{Z}_{j}\rVert^{\beta}.$ (11)
@@
where $\boldsymbol{S}_{t}=\operatorname{diag}(\lambda_{i,t})_{i=1}^{N}$ is a diagonal matrix containing the eigenvalues $\lambda_{i,t}$ of $\boldsymbol{\Sigma}_{t}$, and $\boldsymbol{U}_{t}$ is the orthonormal matrix of corresponding eigenvectors. We then define
@@
where $\mathbf{w}_{t}$ is a random vector with each element following a standard Gaussian distribution, i.e., ${w}_{i,t}\sim\mathcal{N}\left(0,1\right)$. We can then apply Eq. (8) individually to each element and formulate the MVG-CRPS mimicking Eq. (9) for multivariate problem:
$\mathcal{L}\left(\mathcal{N}\left(\boldsymbol{\mu}_{t},\boldsymbol{\Sigma}_{t}\right),\mathbf{z}_{t}\right)=\sum_{i=1}^{N}\sqrt{\lambda_{i,t}}\operatorname{CRPS}\left(\Phi,{w}_{i,t}\right).$ (15)
The overall loss function for training the model is then formulated over an observation period $T$:
$\mathcal{L}=\sum_{t=1}^{T}\mathcal{L}\left(\mathcal{N}\left(\boldsymbol{\mu}_{t},\boldsymbol{\Sigma}_{t}\right),\mathbf{z}_{t}\right).$ (16)
The primary advantage of this loss function is its ability to exploit the closed-form expression of $\operatorname{CRPS}$ by decorrelating time series variables through whitening. Applied to each transformed (whitened) variable, it evaluates the marginal distributions in the transformed space, where the whitening process is determined by the original covariance matrix. Consequently, the optimization process retains information about the covariance structure of the original distribution. Under the Gaussian assumption, MVG-CRPS is a strictly proper scoring rule (see Appendix §A).
@@ 4.1 Datasets and Models
We apply the MVG-CRPS to two multivariate forecasting tasks: multivariate autoregressive forecasting and univariate Seq2Seq forecasting. For the first task, we employ two benchmark models: the RNN-based GPVar (Salinas et al., 2019) and a decoder-only Transformer (Radford et al., 2018). For the second task, we use the N-HiTS model, which is based on multi-layer perceptrons (MLPs) (Challu et al., 2023).
@@
To generate the distribution parameters for probabilistic forecasting, we employ a Gaussian distribution head based on the hidden state $\mathbf{h}_{i,t}$ produced by the model. Specifically, for the multivariate autoregressive forecasting, following Salinas et al. (2019), we parameterize the mean vector as $\boldsymbol{\mu}\left(\mathbf{h}_{t}\right)=\left[\mu_{1}\left(\mathbf{h}_{1,t}\right),\dots,\mu_{N}\left(\mathbf{h}_{N,t}\right)\right]^{\top}\in\mathbb{R}^{N}$ and adopt a low-rank-plus-diagonal parameterization of the covariance matrix $\boldsymbol{\Sigma}\left(\mathbf{h}_{t}\right)=\boldsymbol{L}_{t}\boldsymbol{L}_{t}^{\top}+\operatorname{diag}{(\mathbf{d}_{t})}$, where $\mathbf{d}_{t}=[d_{1}\left(\mathbf{h}_{1,t}\right),\ldots,d_{N}\left(\mathbf{h}_{N,t}\right)]^{\top}\in\mathbb{R}_{+}^{N}$ and $\boldsymbol{L}_{t}=\left[\mathbf{l}_{1}\left(\mathbf{h}_{1,t}\right),\dots,\mathbf{l}_{N}\left(\mathbf{h}_{N,t}\right)\right]^{\top}\in\mathbb{R}^{N\times R}$, $R\ll N$ is the rank parameter. Here, $\mu_{i}(\cdot)$, $d_{i}(\cdot)$, and $\mathbf{l}_{i}(\cdot)$ are the mapping functions that generate the mean and covariance parameters for each time series $i$ based on the hidden state $\mathbf{h}_{i={1:N},t}$. In practice, we use shared mapping functions across all time series, denoted as $\mu_{i}=\tilde{\mu}$, $d_{i}=\tilde{d}$, and $\mathbf{l}_{i}=\tilde{\mathbf{l}}$. This parameterization ensures that $\boldsymbol{\Sigma}(\mathbf{h}_{t})$ is positive definite and efficiently parameterized. The diagonal component provides stability, while the low-rank component captures the covariance structure. The Gaussian assumption also enables the use of random subsets of time series (i.e., batch size $B\leq N$) for model optimization in each iteration, making it feasible to apply our method to high-dimensional time series datasets. Similarly, in the univariate Seq2Seq forecasting task, the mean $\boldsymbol{\mu}\left(\mathbf{h}_{i}\right)$ and covariance $\boldsymbol{\Sigma}(\mathbf{h}_{i})$ are defined over the forecast horizon for each specific time series, based on the hidden states $\mathbf{h}_{i,t=T+1:T+Q}$. As a result, we can model the joint distribution $p\left(\mathbf{z}_{i,{T+1}:{T+Q}}\right)$ over the forecasted values.
@@
We implemented our models using PyTorch Forecasting (Beitner, 2020), with input data consisting of lagged time series values and covariates. We conducted extensive experiments on diverse real-world time series datasets obtained from GluonTS (Alexandrov et al., 2020) (see Appendix §B). The prediction horizon ($Q$) and the number of rolling evaluations were taken from the configuration within GluonTS, where we follow the default setting by equating the context range to the prediction range, i.e., $P=Q$. Each dataset was sequentially split into training, validation, and testing sets. Each time series was individually normalized using a scaler fitted to its own training data (Salinas et al., 2020; Kim et al., 2021). Predictions were then rescaled to their original values for computing evaluation metrics. For comprehensive details on the experimental setup, we refer readers to Appendix §C.
4.2 Toy Example
We first perform a toy experiment following Roordink & Hess (2023) using a true distribution $P=\mathcal{N}\left(\begin{bmatrix}1\\ -1\end{bmatrix},\begin{bmatrix}1&0.8\\ 0.8&4\end{bmatrix}\right)$ and a predictive distribution $Q=\mathcal{N}\left(\begin{bmatrix}\mu\\ -1\end{bmatrix},\begin{bmatrix}\sigma^{2}&2\rho\sigma\\ 2\rho\sigma&4\end{bmatrix}\right)$, where we control the deviation of the three parameters $\mu,\rho,\sigma$ to study the various properties of different scores. As shown in Fig. 3, the log-score exhibits high sensitivity to parameter deviations, particularly to inaccuracies in standard deviation ($\sigma$) and correlation coefficient ($\rho$) estimation. While the energy score is less sensitive than the log-score, it can occasionally produce slightly inconsistent scores due to its sample-based approximation (ideally, larger deviations result in higher scores). Furthermore, the energy score demonstrates lower sensitivity to deviations in correlation. In contrast, the proposed MVG-CRPS provides a well-balanced sensitivity across all three parameters while maintaining consistent scoring patterns, owing to its closed-form expression. Additionally, the strict properness of MVG-CRPS under the Gaussian assumption implies that the score is uniquely minimized when $P=Q$.
Figure 3: Sensitivity of scoring rules to parameter deviations in the predicted mean, standard deviation, and correlation coefficient from the true data distribution ($\mu_{\text{true}}=1,\sigma_{\text{true}}=1,\rho_{\text{true}}=0.4)$. The energy score is computed with a sample size of 500.
@@ 4.3 Quantitative Evaluation
We evaluate the MVG-CRPS by comparing it with models trained using the log-score and energy score. The evaluation metrics are $\operatorname{CRPS}_{\text{sum}}$ and energy score (Eq. (11)). $\operatorname{CRPS}_{\text{sum}}$ is computed by summing both the forecast and ground-truth values across all time series and then calculating the $\operatorname{CRPS}$ (Eq. (7)) over the resulting sums (Salinas et al., 2019; Drouin et al., 2022; Ashok et al., 2024):
$\operatorname{CRPS}_{\text{sum}}=\operatorname{\mathbb{E}}_{t}\left[\operatorname{CRPS}\left(F_{t},\sum_{i}z_{i,t}\right)\right],$ (17)
where the empirical $F_{t}$ is calculated by aggregating samples across time series, with 100 samples drawn to compute $\operatorname{CRPS}_{\text{sum}}$. The evaluation using the energy score is provided in the Appendix §D.
Table 1: Comparison of $\operatorname{CRPS}_{\text{sum}}$ across different scoring rules in the multivariate autoregressive forecasting task. The best scores are in boldface. MVG-CRPS scores are underlined when they are not the best overall but exceed the log-score. Mean and standard deviation are reported from 10 runs of each model.
@@
Table 2: Comparison of training cost (in minutes) for GPVar under different scoring rules in the multivariate autoregressive forecasting task.
@@
Table 1 presents a comparison of $\operatorname{CRPS}_{\text{sum}}$ across different scoring rules in the multivariate autoregressive forecasting task, underscoring the effectiveness of the MVG-CRPS. MVG-CRPS consistently outperforms the log-score across most datasets. For instance, on the $\mathtt{electricity}$ and $\mathtt{tourism}$ datasets, MVG-CRPS achieves scores of 0.0249 and 0.2004, respectively, outperforming the scores of 0.0419 and 0.2217 achieved by the log-score. This indicates that MVG-CRPS leads to models that produce higher-quality forecasts. As discussed in later sections, this improvement is attributed to MVG-CRPS being less sensitive to outliers.
@@ 4.4.1 Multivariate Autoregressive Forecasting
Figure 4: Comparison of output covariance matrices $\boldsymbol{\Sigma}_{t}$ from GPVar on the $\mathtt{elec\_weekly}$ dataset. The top and bottom rows display covariance matrices from models trained using the log-score and MVG-CRPS. For visual clarity, covariance values are clipped between 0 and 0.6.
@@
Figure 5: Comparison of probabilistic forecasts from GPVar on the $\mathtt{electricity}$ dataset. The first and second rows display forecasts from models trained using the log-score and MVG-CRPS.
@@ 4.4.2 Univariate Seq2Seq Forecasting
Figure 6: Comparison of output covariance matrices $\boldsymbol{\Sigma}_{i}$ from N-HiTS on the $\mathtt{traffic}$ dataset. For visual clarity, covariance values are clipped between 0 and 1.0.
Figure 7: Comparison of probabilistic forecasts from N-HiTS on the $\mathtt{m4\_hourly}$ dataset.
@@
Another example is shown in Fig. 7, which depicts the probabilistic forecasts generated by the N-HiTS model on the $\mathtt{m4\_hourly}$ dataset. The MVG-CRPS-trained model produces narrower prediction intervals, suggesting greater confidence and improved calibration compared to the log-score-trained model. In contrast, the log-score-trained model generates overly wide intervals for simple time series with strong cyclical patterns. Additionally, the MVG-CRPS-trained model achieves higher accuracy for longer-term forecasts.
These qualitative observations highlight the benefits of MVG-CRPS for probabilistic forecasting. By reducing sensitivity to outliers and providing more stable covariance estimates, MVG-CRPS supports models in delivering accurate and reliable forecasts.